OpenStudy (kpop4life123):

A line passes through (2, 8) and (4, 12). Which equation best represents the line? (4 points) y = 1/2 x + 6 y = 2x + 7 y = 2x + 4 y = 1/2 x + 10

OpenStudy (anonymous):

do samething as before replace numbers within the equation

OpenStudy (anonymous):

such grammar .-. same thing *

OpenStudy (kpop4life123):

is the 2 and 4 the yaxis?

OpenStudy (kpop4life123):

sorry im not good in math

OpenStudy (anonymous):

the 2 and 4 are the x-axis

OpenStudy (kpop4life123):

oh ok thanks :)

OpenStudy (anonymous):

remember its (x,y)

OpenStudy (lynfran):

you need to first find the slope for the line using the 2 given points

OpenStudy (lynfran):

slope= y/x so, (8-12) / (2-4)=??

OpenStudy (kpop4life123):

its c right

OpenStudy (lynfran):

yes

OpenStudy (lynfran):

slope=2 so y=2x+b using point (2,8) (8)=2(2)+b 8-4=b 4=b so y=2x+4

OpenStudy (kpop4life123):

thanks both of yall :)